OpenStudy (anonymous):

How many atoms are in 15 grams of Carbon?

OpenStudy (anonymous):

No of atoms = Na * No of Moles * Atoms

OpenStudy (anonymous):

Moles = Given mass/MM

OpenStudy (anonymous):

Na = 6.022 * 10 ^23

OpenStudy (anonymous):

Atoms = 1 Since u r luking for C if the compound is SO2 then it is 2 + 1 = 3

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

is the answer 1.25 x 10^23

OpenStudy (anonymous):

1.25 * 6.022 * 10^23

OpenStudy (anonymous):

so 7.53 x 10^23?